# On the stability, complexity, and distribution of similarity classes of the longest edge bisection process for triangles

### v2

@@ Abstract.
The Longest Edge Bisection of a triangle is performed by joining the midpoint of its longest edge to the opposite vertex. Applying this procedure iteratively produces an infinite family of triangles. Surprisingly, a classical result of Stynes (1980) shows that for any initial triangle, this infinite family falls into finitely many similarity classes.
While the set of classes is finite, it turns out that a far smaller, stable subset of “fat” triangles effectively dominates the final mesh structure. We prove the following asymptotic area distribution result: for every initial triangle, the portion of area occupied by terminal quadruples tends to one, with the convergence occurring at an exponential rate. In fact, we provide the precise distribution of triangles in every step. We introduce the bisection graph and use spectral methods to establish this result.
Given this dominance, we provide a complete characterization of triangles possessing a single terminal quadruple, while conversely exhibiting a sequence of triangles with an unbounded number of terminal quadruples. Furthermore, we reveal several fundamental geometric properties of the points of a terminal quadruple, laying the groundwork for studying the geometric distribution of the entire orbit. Our analysis leverages the hyperbolic geometry framework of Perdomo and Plaza (2014) and refines their techniques.
@@ 1.1. Background
The Longest Edge Bisection (LEB) process is a straightforward iterative procedure formally initiated by Rosenberg and Stenger [RS75]. In each step, the median to the longest edge is added, splitting the triangle into two daughter triangles. Rosenberg and Stenger established the critical property of non-degeneracy, which guarantees geometric quality. They proved that for an initial triangle $\Delta_{0}$ with smallest angle $\alpha_{0}$, every angle generated by the infinite process is bounded from below by $\frac{\alpha_{0}}{2}$. Additionally, a useful triangulation process requires the convergence of the mesh diameter to zero. Kearfott [KEA78] first proved this for LEB in arbitrary dimensions; later, Stynes [STY79, STY80] and Adler [ADL83] provided sharpened bounds on the convergence rates.
A landmark result, known as the Finite Similarity Classes (FSC) Theorem, was established by Stynes [STY80, Corollary 1]. This theorem asserts that the infinite family of triangles generated by iteratively bisecting a single initial triangle falls into a finite number of similarity classes. This result was later reproved by Adler [ADL83] using an elegant inductive argument, and subsequently by Perdomo and Plaza [PP14], who recast the analysis within the framework of hyperbolic geometry, a setting that proved to be particularly natural.
In their paper, Rosenberg and Stenger identified a specific region where triangles enter a periodic orbit of length $4$ under the $\mathrm{LEB}$ process [RS75, Lemma 4]; we refer to these orbits as terminal sets. Building on this work and the FSC Theorem, Stynes [STY80, Corollary 2] demonstrated that the proportion of the area occupied by triangles in these terminal sets converges to $1$.
These results pave the way for a deeper, quantitative analysis of the $\mathrm{LEB}$ process as a dynamical system. While it is known that the variety of shapes is finite and that terminal sets asymptotically exhaust the area of the initial triangle, this does not describe the rate of convergence or the precise distribution of triangle shapes. Several fundamental questions naturally arise:
At what rate does the total area of the mesh become dominated by these stable shapes? What are the specific geometric properties of these terminal sets? Furthermore, how does the number of distinct terminal sets depend on the geometry of the initial triangle? Answering these questions is essential for a comprehensive understanding of the stability and complexity of the $\mathrm{LEB}$ process. We provide a precise formulation of these problems in §1.2 and address them as the main contribution of this paper. Partial results in these directions were previously noted in [GGR07].
@@ 1.2. Preliminaries and main results
The 2014 paper [PP14] by Perdomo and Plaza provided a new proof of the FSC theorem, employing the machinery of hyperbolic dynamics in a moduli space of triangle shapes, modeled in the hyperbolic upper half-plane $\mathbb{H}$, to analyze the bisection process as a dynamical system. Given a triangle $\Delta_{0}$, by
  1. (1)
    scaling the edge lengths so that the longest edge has length $1$,
  2. (2)
    mapping the longest edge to the interval $[0,1]$ on the real axis, and
  3. (3)
    placing the third vertex $z$ opposite to the longest edge, on the upper half plane, and so that the shortest edge is on the left,
we bijectively identify every similarity class of triangles with a unique point $z$ in the domain $D\subseteq\mathbb{H}$ defined by (see Figure 1).
$D=\left\{z\in\mathbb{C}\ \middle|\ \mathrm{Re}(z)\leq\frac{1}{2},\>\mathrm{Im}(z)>0,\>\left|{z-1}\right|\leq 1\right\}.$
Relying on this identification, we refer to a point $z\in D$ as a triangle. For such $z\in D$, the bisection by the median to the longest edge defines two function $L,R:D\to D$, where $L(z)$ is the left triangle and $R(z)$ is the right triangle. As explained in [PP14], the region $D$ is subdivided into $6$ subregions by the geodesics $\mathrm{Re}(z)=\frac{1}{4}$, $\left|{z}\right|=\frac{1}{2}$ and $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$ (see Figure 1), and the functions $L$ and $R$ are defined in a piecewise manner, by a Möbius or anti-Möbius transformation on each one of these regions. The formulas and the geometric interpretation of these transformations are presented in Table 1. An illustrative application presenting the geometry of the various transformations is available at https://danielkalmanovich.wixsite.com/website/leb. The reader is referred to [AHL78, §3] for more details on the action by Möbius transformations on the upper half-plane.
@@
(b) The subdivision of $D$ into $6$ subregions
Figure 1. The normalized region $D$ and its subdivision.
Region Function Geometric interpretation in $D$
@@
II $L(z)=\frac{-1}{2z-1}$ An elliptic Möbius map with a fixed point at $\frac{1}{4}+\frac{\sqrt{(}7)}{4}i\in D$. Points rotate counterclockwise by an angle of $\arccos\left(-\frac{3}{4}\right)$ about the fixed point.
@@
IV $L(z)=\frac{2z-1}{2z}$ An elliptic Möbius map with a fixed point at $\frac{1}{2}+\frac{1}{2}i\in D$. Points rotate clockwise by an angle of $\frac{\pi}{2}$ about the fixed point.
@@
$\text{I}\cup\text{II}\cup\text{III}$ $R(z)=\frac{-1}{2z-2}$ An elliptic Möbius map with a fixed point at $\frac{1}{2}+\frac{1}{2}i\in D$. Points rotate counterclockwise by an angle of $\frac{\pi}{2}$ about the fixed point.
@@
As we will show, an exceptional property of the dynamics of the $\mathrm{LEB}$ process is that, regardless of the initial triangle, the process eventually produces predominantly triangles from a certain subregion of $\text{I}\cup\text{IV}$, with notable geometric properties. Furthermore, the orbit of such a triangle is periodic of size $4$, thus we call it a terminal quadruple. These facts immediately raise several questions that motivate this paper.
@@
    How many distinct terminal quadruples can arise from a given initial triangle?
@@
To answer these questions, we introduce the bisection graph (see §3) of a triangle, and use spectral methods to analyze it. Furthermore, we adopt and refine the methods used in [PP14] to study the geometric properties of the orbit of $z$ under iterative $\mathrm{LEB}$. We denote by $\mathrm{LEB}(z)$ the orbit of $z$, by $q(z)$ the number of terminal quadruples that arise from $z$, and by $l(z)=\left|{\mathrm{LEB}(z)}\right|$. Our main results are the following.
@@
For a triangle $z$ and $j\in\mathbb{N}$ let $w_{j}\in\mathbb{R}^{l(z)}$ be the probability vector that describes the partition of the area of $z$ into triangles of different similarity classes, after $j$ steps. Then, for every triangle $z$ we have the following:
@@
    The limits $w_{even}=\displaystyle{\lim_{j\to\infty}w_{2j}}$ and $w_{odd}=\displaystyle{\lim_{j\to\infty}w_{2j+1}}$ exist, and the non-zero entries of $w_{even}$ and $w_{odd}$ are those corresponding to elements in terminal quadruples.
@@
    There exists $\xi\in(0,1)$ such that for every $j\in\mathbb{N}$, after $j$ steps, the area of $z$ that is occupied by triangles that do not belong to terminal quadruples is $O(\xi^{j})$.
This theorem immediately implies Stynes’ result, stating that the area of any initial triangle $z$ gets filled with triangles that belong to terminal quadruples (see [STY80, Corollary 2]). Furthermore, Theorem 1 exhibits the precise distribution vector at every step and the rate of convergence.
It follows that the number of triangle shapes that one sees is essentially $4q(z)$ and not $l(z)$. For instance, taking an acute isosceles triangle with apex angle $\alpha$ we have $l(z)=\Theta\left(\log\left(\frac{1}{\alpha}\right)\right)$ [GGR07, Lemma 2], while according to Theorem 2 below we have $q(z)=1$. In general, the following theorems shed light on the quantity $q(z)$, for $z$ in the different regions.
@@
For every $n\in\mathbb{N}$ there are triangles $z$ in regions V and VI for which $q(z)>n$.
@@ 2. Terminal quadruples
We start this section with a classification of the non-generic triangles in $D$, followed by a fundamental geometric property of terminal quadruples. In the last two subsections we prove Theorem 2 and Theorem 3.
@@ 2.1. Generic triangles and triangles with orbits of size less than 44
We begin with the definition of a generic triangle with respect to the $\mathrm{LEB}$ process. It can happen that $L(z)$ is similar to $R(z)$, or that one of them is similar to $z$, thus we define:
@@
A triangle $z\in D$ is generic if for each $w\in\mathrm{LEB}(z)$, the points $w,L(w)$ and $R(w)$ are all distinct.
@@
A triangle $z\in D$ is generic if and only if $\mathrm{LEB}(z)$ does not intersect the following $3$ geodesics:
@@
First, we address points $w\in D$ that are fixed under either $L$ or $R$. Since $L_{\text{I}}$ is an inversion with respect to the geodesic $\left|{w}\right|=\frac{\sqrt{2}}{2}$, we have $L_{\text{I}}(w)=w$ for points lying on this geodesic. Similarly, $R_{\text{IV}\cup\text{V}\cup\text{VI}}$ is an inversion with respect to the geodesic $\left|{w-1}\right|=\frac{\sqrt{2}}{2}$ and so $R_{\text{IV}\cup\text{V}\cup\text{VI}}(w)=w$ there. Each one of $L_{\text{II}},L_{\text{IV}}$ and $R_{\text{I}\cup\text{II}\cup\text{III}}$ has a unique fixed point, which is on one of these two geodesics, and $L_{\text{III}},L_{\text{V}}$ and $L_{\text{VI}}$ have no fixed points in $D$. Thus $w=R(w)$ or $w=L(w)$ if and only if $\left|{w-1}\right|=\frac{\sqrt{2}}{2}$ or $\left|{w}\right|=\frac{\sqrt{2}}{2}$.
Next, we show that $L(w)=R(w)$ if and only if $\mathrm{Re}(w)=\frac{1}{2}$. Notice that for every $w\in D$, the triangle $R(w)$ is not acute. In case $L(w)$ is obtuse, one can easily see that $L(w)\neq R(w)$ because the obtuse angle of $L(w)$ is strictly smaller than that of $R(w)$. So we may assume that $L(w)$ is not obtuse. Thus $L(w)=R(w)$ if and only if $L(w)$ and $R(w)$ are the same right-angled triangle. Observe that if $L(w)=R(w)$ then $w$ is isosceles, and if the apex of an isosceles triangle $w$ is $\frac{\pi}{3}$, then its longest edge is its leg, thus $L(w)\neq R(w)$ in that case. Hence, we conclude that $L(w)=R(w)$ if and only if $w$ is an isosceles triangle with apex $\geq\frac{\pi}{3}$. Thus $L(w)=R(w)$ if and only if $\mathrm{Re}(w)=\frac{1}{2}$. ∎
@@
(a) The non-generic triangles in $D$: The triangles on the upper red curve satisfy $z=L(z)$, the triangles on the lower red curve satisfy $z=R(z)$. The triangles on the green curve satisfy $L(z)=R(z)$.
@@
(b) The non-generic triangles in the shaded terminal region AA: the blue $\zeta$ generates 11 triangle. The triangles on the red curves generate 22 triangles. The triangles on the green curves generate 33 triangles.
@@
In view of Proposition 5, the set of non-generic triangles is contained in a countable union of arcs in $D$, i.e., the inverse images of the arcs described in (1) with respect to iterative applications of $L$ and $R$, thus has measure zero. In particular, the set of generic triangles is of full measure.
Since any region I triangle generates at most $4$ similarity classes and I is invariant under $L$, it follows that for any triangle $z$ in the region
$A=\text{I}\cup R(\text{I})\subseteq\text{I}\cup\text{IV}$
we have $l(z)\leq 4$. It is easy to verify that for any triangle $z\in D\setminus A$ we have $l(z)\geq 5$, thus $A$ is precisely the locus $\left\{z\in D\ \middle|\ l(z)\leq 4\right\}$. We refer to $A$ as the terminal region and the orbit $\mathrm{LEB}(z)$ of any $z\in A$ is called a terminal quadruple. This terminology applies consistently to all $z\in A$: generic triangles produce orbits of $4$ distinct elements, while non-generic triangles produce orbits of $4$ elements, some of which are repeated. The precise characterization of these non-generic cases is provided in the following corollary.
@@
Let $z\in D$ and set $\zeta=\frac{1}{2}+\frac{1}{2}i$, then:
@@
    $l(z)=2$ if and only if $z\in\left\{w\in A\ \middle|\ \left|{w}\right|=\frac{\sqrt{2}}{2}~~\text{ or }~~|w-1|=\frac{\sqrt{2}}{2}\right\}\setminus\{\zeta\}$.
@@
    $l(z)=3$ if and only if $z\in\left\{w\in A\ \middle|\ \left|{w-\frac{1}{2}}\right|=\frac{1}{2}~~\text{ or }~~\mathrm{Re}(w)=\frac{1}{2}\right\}\setminus\{\zeta\}$.
@@
Since the geodesic $\left|{w-\frac{1}{2}}\right|=\frac{1}{2}$ in $A$ is the image of the geodesic $\mathrm{Re}(w)=\frac{1}{2}$ in AA under $L$ and $R$, this is a direct consequence of Proposition 5 (see Figure 2). ∎

### v3

@@ Abstract.
The Longest Edge Bisection of a triangle is performed by joining the midpoint of its longest edge to the opposite vertex. Applying this procedure iteratively produces an infinite family of triangles. Surprisingly, a classical result of Stynes (1980) shows that for any initial triangle, the elements of this infinite family fall into finitely many similarity classes.
While the set of classes is finite, it turns out that a far smaller, periodic subset of “fat” triangles effectively dominates the final mesh structure. This subset is comprised of periodic orbits of length four, which we refer to as terminal quadruples. We prove the following asymptotic area distribution result: for every initial triangle, the portion of area occupied by these terminal quadruples tends to one, with the convergence occurring at an exponential rate. In fact, we provide the precise distribution of triangles in every step. We introduce the bisection graph and use spectral methods to prove this result.
Given this dominance, we provide a complete characterization of triangles possessing a single terminal quadruple, while conversely exhibiting a sequence of triangles with an unbounded number of terminal quadruples. Furthermore, we reveal several fundamental geometric properties of the points of a terminal quadruple, laying the groundwork for studying the geometric distribution of the entire orbit.
@@ 1.1. Background
The Longest Edge Bisection ($\mathrm{LEB}$) process is a straightforward iterative procedure, formally defined and studied by Rosenberg and Stenger in 1975 [RS75]. In each step, every existing triangle is bisected by the median to its own longest edge, splitting it into two daughter triangles. Rosenberg and Stenger proved the critical property of non-degeneracy: for an initial triangle $\Delta_{0}$ with smallest angle $\alpha_{0}$, the angles of all triangles generated by the infinite process are bounded from below by $\frac{\alpha_{0}}{2}$. Another requirement for a useful triangulation process is the convergence of the mesh diameter to zero. Kearfott [KEA78] first proved this for $\mathrm{LEB}$ in arbitrary dimensions; later, Stynes [STY79, STY80] and Adler [ADL83] provided sharpened bounds on the convergence rates.
A landmark result, known as the Finite Similarity Classes (FSC) theorem, was established by Stynes [STY80, Corollary 1]. This theorem states that, although iteratively bisecting a single initial triangle produces infinitely many triangles, they belong to only finitely many similarity classes. This result was later reproved by Adler [ADL83] using an elegant inductive argument, and subsequently by Perdomo and Plaza [PP14], who recast the analysis within the framework of hyperbolic geometry, a setting that proved to be particularly natural.
In their paper, Rosenberg and Stenger identified a specific region where triangles enter a periodic orbit of length $4$ under the $\mathrm{LEB}$ process [RS75, Lemma 4], forming what we call a terminal quadruple. Building on this work and the FSC theorem, Stynes [STY80, Corollary 2] demonstrated that the proportion of the area occupied by triangles in these terminal quadruples converges to $1$.
These results pave the way for a deeper, quantitative analysis of the $\mathrm{LEB}$ process as a dynamical system. While it is known that the variety of shapes is finite and that terminal quadruples asymptotically exhaust the area of the initial triangle, this does not describe the rate of convergence or the precise distribution of triangle shapes. Partial results in these directions were previously noted in [GGR07].
@@ 1.2. Preliminaries and main results
To analyze the $\mathrm{LEB}$ process as a dynamical system, we employ the framework introduced by Perdomo and Plaza [PP14], which utilizes the machinery of hyperbolic dynamics in a moduli space of triangle shapes modeled in the hyperbolic upper half-plane $\mathbb{H}$. Given an arbitrary triangle, we normalize it by scaling its longest edge to length 1 and mapping this edge to the interval $[0,1]$ on the real axis. By placing the third vertex in the upper half-plane $\mathbb{H}$ such that the shortest edge lies on the left, we bijectively identify every similarity class of triangles with a unique point $z$ in the domain $\mathcal{D}\subseteq\mathbb{H}$ defined by (see Figure 1):
$\mathcal{D}=\left\{z\in\mathbb{C}\ \middle|\ \mathrm{Re}(z)\leq\frac{1}{2},\>\mathrm{Im}(z)>0,\>\left|{z-1}\right|\leq 1\right\}.$
In light of this bijection, we will refer to elements $z\in\mathcal{D}$ interchangeably as points or triangles. For such $z\in\mathcal{D}$, bisecting the triangle by the median to its longest edge defines two maps $L,R:\mathcal{D}\to\mathcal{D}$, where $L(z)$ is the left triangle and $R(z)$ is the right triangle. As explained in [PP14], the domain $\mathcal{D}$ is subdivided into $6$ subregions by the geodesics $\mathrm{Re}(z)=\frac{1}{4}$, $\left|{z}\right|=\frac{1}{2}$ and $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$ (see Figure 1), where the functions $L$ and $R$ are defined in a piecewise manner, by a Möbius or anti-Möbius transformation on each one of these subregions. We consider the action of the semigroup $\langle{L},{R}\rangle$ on $\mathcal{D}$ and refer to words in this semigroup as the functions that they correspond to. We also adopt standard terminology from dynamical systems, such as the orbit of a point. The explicit formulas and geometric interpretations of these transformations are summarized in Table 1. An illustrative application presenting the geometry of the various transformations is available at https://danielkalmanovich.wixsite.com/website/leb. The reader is referred to [AHL78, §3] for more details on the action by Möbius transformations on the upper half-plane.
@@
(b) The subdivision of $\mathcal{D}$ into $6$ subregions
Figure 1. The normalized region $\mathcal{D}$ and its subdivision.
Region Function Geometric interpretation in $\mathcal{D}$
@@
II $L(z)=\frac{-1}{2z-1}$ An elliptic Möbius map with a fixed point at $\frac{1}{4}+\frac{\sqrt{(}7)}{4}i\in\mathcal{D}$. Points rotate counterclockwise by an angle of $\arccos\left(-\frac{3}{4}\right)$ about the fixed point.
@@
IV $L(z)=\frac{2z-1}{2z}$ An elliptic Möbius map with a fixed point at $\frac{1}{2}+\frac{1}{2}i\in\mathcal{D}$. Points rotate clockwise by an angle of $\frac{\pi}{2}$ about the fixed point.
@@
$\text{I}\cup\text{II}\cup\text{III}$ $R(z)=\frac{-1}{2z-2}$ An elliptic Möbius map with a fixed point at $\frac{1}{2}+\frac{1}{2}i\in\mathcal{D}$. Points rotate counterclockwise by an angle of $\frac{\pi}{2}$ about the fixed point.
@@
In view of the fact that the $\mathrm{LEB}$ process eventually produces predominantly terminal quadruples, it seems beneficial to move beyond this qualitative result and achieve a comprehensive quantitative and geometric understanding of the process. We are led to the following questions:
@@
    How many distinct terminal quadruples can arise from a given initial triangle $z$?
@@
To answer these questions, we introduce the bisection graph of a triangle (see §3) and analyze it using spectral methods. Furthermore, we adopt and refine the methods used in [PP14] to study the geometric properties of the orbit of $z$ under iterative $\mathrm{LEB}$. We denote by:
  • $\mathrm{LEB}(z)$ the orbit of $z$,
  • $q(z)$ the number of terminal quadruples that arise from $z$,
  • $l(z)=\left|{\mathrm{LEB}(z)}\right|$.
Our main results are the following.
@@
For a triangle $z$ and $j\in\mathbb{N}$, let $w_{j}\in\mathbb{R}^{l(z)}$ be the probability vector describing the partition of the area of $z$ into triangles of different similarity classes, after $j$ bisection steps. Then for every triangle $z$ we have:
@@
    The limits $w_{even}=\displaystyle{\lim_{j\to\infty}w_{2j}}$ and $w_{odd}=\displaystyle{\lim_{j\to\infty}w_{2j+1}}$ exist, and their non-zero entries correspond to elements in terminal quadruples.
@@
    There exists $\xi\in(0,1)$ such that for every $j\in\mathbb{N}$, the area of $z$ that is occupied by triangles not belonging to terminal quadruples after $j$ steps is $O(\xi^{j})$.
This theorem immediately implies Stynes’ result, stating that the area of any initial triangle $z$ gets filled with triangles that belong to terminal quadruples (see [STY80, Corollary 2]). Furthermore, Theorem 1 provides the precise distribution vector at every step and the rate of convergence.
It follows that the number of triangle shapes that one sees is essentially $4q(z)$ rather than $l(z)$. For instance, taking an acute isosceles triangle with apex angle $\alpha$ we have $l(z)=\Theta\left(\log\left(\frac{1}{\alpha}\right)\right)$ (see  [GGR07, Lemma 2] for the upper bound, and the lower bound is obtained by inspecting the function $L$ in region III), while according to Theorem 2 below we have $q(z)=1$. In general, the following theorems shed light on the quantity $q(z)$, for $z$ in the different regions.
@@
For every $n\in\mathbb{N}$, there are triangles $z$ in regions V and VI for which $q(z)>n$.
@@ 2. Terminal quadruples
We start this section with the definition of the generic triangles with respect to the $\mathrm{LEB}$ process and their classification.
@@ 2.1. Generic triangles and triangles with orbits of size less than 44
It can happen that $L(z)$ is similar to $R(z)$, or that one of them is similar to $z$, thus we define:
@@
A triangle $z\in\mathcal{D}$ is generic if for each $w\in\mathrm{LEB}(z)$, the points $w,L(w)$ and $R(w)$ are all distinct.
@@
A triangle $z\in\mathcal{D}$ is generic if and only if $\mathrm{LEB}(z)$ does not intersect the following $3$ geodesics:
@@
First, we address points $w\in\mathcal{D}$ that are fixed under either $L$ or $R$. Since $L_{\text{I}}$ is an inversion with respect to the geodesic $\left|{w}\right|=\frac{\sqrt{2}}{2}$, we have $L_{\text{I}}(w)=w$ for points lying on this geodesic. Similarly, $R_{\text{IV}\cup\text{V}\cup\text{VI}}$ is an inversion with respect to the geodesic $\left|{w-1}\right|=\frac{\sqrt{2}}{2}$ and so $R_{\text{IV}\cup\text{V}\cup\text{VI}}(w)=w$ there. Each one of $L_{\text{II}},L_{\text{IV}}$ and $R_{\text{I}\cup\text{II}\cup\text{III}}$ has a unique fixed point, which is on one of these two geodesics, and $L_{\text{III}},L_{\text{V}}$ and $L_{\text{VI}}$ have no fixed points in $\mathcal{D}$. Thus $w=R(w)$ or $w=L(w)$ if and only if $\left|{w-1}\right|=\frac{\sqrt{2}}{2}$ or $\left|{w}\right|=\frac{\sqrt{2}}{2}$.
Next, we show that $L(w)=R(w)$ if and only if $\mathrm{Re}(w)=\frac{1}{2}$. Notice that for every $w\in\mathcal{D}$, the triangle $R(w)$ is not acute. In case $L(w)$ is obtuse, one can easily see that $L(w)\neq R(w)$ because the obtuse angle of $L(w)$ is strictly smaller than that of $R(w)$. So we may assume that $L(w)$ is not obtuse. Thus $L(w)=R(w)$ if and only if $L(w)$ and $R(w)$ are the same right-angled triangle. Observe that if $L(w)=R(w)$ then $w$ is isosceles, and if the apex of an isosceles triangle $w$ is $\frac{\pi}{3}$, then its longest edge is its leg, thus $L(w)\neq R(w)$ in that case. Hence, we conclude that $L(w)=R(w)$ if and only if $w$ is an isosceles triangle with apex $\geq\frac{\pi}{3}$. Thus $L(w)=R(w)$ if and only if $\mathrm{Re}(w)=\frac{1}{2}$. ∎
@@
(a) The non-generic triangles in $\mathcal{D}$: The triangles on the upper red curve satisfy $z=L(z)$, the triangles on the lower red curve satisfy $z=R(z)$. The triangles on the green curve satisfy $L(z)=R(z)$.
@@
(b) The non-generic triangles in the shaded terminal region $\mathcal{A}$: the blue $\zeta$ generates 11 triangle. The triangles on the red curves generate 22 triangles. The triangles on the green curves generate 33 triangles.
@@
In view of Proposition 5, the set of non-generic triangles is contained in a countable union of arcs in $\mathcal{D}$, i.e., the inverse images of the arcs described in (1) with respect to iterative applications of $L$ and $R$, thus has measure zero. In particular, the set of generic triangles is of full measure.
It is well-known that any region I triangle generates at most $4$ similarity classes (see e.g. [PP14]). Generic triangles produce orbits of $4$ distinct elements, while non-generic triangles produce orbits of $4$ elements, some of which are repeated. Moreover, this orbit exhibits a specific periodic structure under the bisection maps, as in the following definition.
Definition 7.
An orbit $\{z_{1},z_{2},z_{3},z_{4}\}$ (with not necessarily distinct elements) under the $\mathrm{LEB}$ process is called a terminal quadruple if $L$ and $R$ map $\{z_{1},z_{4}\}$ to $\{z_{2},z_{3}\}$, and vice versa.
In fact, the locus of triangles exhibiting this behavior is larger. Since I is invariant under $L$, it follows that any triangle $z$ in the region
$\mathcal{A}=\text{I}\cup R(\text{I})\subseteq\text{I}\cup\text{IV}$
is such a triangle (see [RS75, Lemma 4], where this was already established). The fact that $\mathcal{A}$ is precisely the locus of terminal quadruples is shown in Corollary 12 below.
The precise characterization of the non-generic terminal quadruple triangles is provided in the following corollary.
@@
Let $z\in\mathcal{D}$ and set $\zeta=\frac{1}{2}+\frac{1}{2}i$, then:
@@
    $l(z)=2$ if and only if $z\in\left\{w\in\mathcal{A}\ \middle|\ \left|{w}\right|=\frac{\sqrt{2}}{2}~~\text{ or }~~|w-1|=\frac{\sqrt{2}}{2}\right\}\setminus\{\zeta\}$.
@@
    $l(z)=3$ if and only if $z\in\left\{w\in\mathcal{A}\ \middle|\ \left|{w-\frac{1}{2}}\right|=\frac{1}{2}~~\text{ or }~~\mathrm{Re}(w)=\frac{1}{2}\right\}\setminus\{\zeta\}$.
@@
Since the geodesic $\left|{w-\frac{1}{2}}\right|=\frac{1}{2}$ in $\mathcal{A}$ is the image of the geodesic $\mathrm{Re}(w)=\frac{1}{2}$ in $\mathcal{A}$ under $L$ and $R$, this is a direct consequence of Proposition 5 (see Figure 2). ∎